A horizontal circular disc with radius r = 0.9681m rotates around its own axis. How long does it take for the disc to spin 1.34 revolutions if the angular velocity is w(t) = 3.343t rad/s?
1
Expert's answer
2021-05-31T17:17:30-0400

"\varphi=\frac{\varepsilon t^2}{2},\\t=\sqrt{\frac{2\varphi}{\varepsilon}}=\sqrt{\frac{4\pi n}{\varepsilon}},"

"t=\sqrt{\frac{4\cdot 3.14\cdot 1.34}{3.343}}=2.2~s."